OpenStudy (adamk):

Find a function for which f(x+3) is equivalent to f(x) + f(3).

OpenStudy (adamk):

I've found a function that is NOT equivalent to f(x) + f(3). f(x) = (x-1) (x-3) (x+2) but I don't know how to find one that is equivalent.

jimthompson5910 (jim_thompson5910):

Claim: If f(x) = m*x, where m is any real number, then f(x+k) = f(x)+f(k) Proof of that claim f(x) = m*x f(x+k) = m*(x+k) f(x+k) = m*x+m*k f(x) = m*x f(k) = m*k f(x)+f(k) = m*x + m*k which means f(x+k) = f(x)+f(k) turns into m*x+m*k = m*x + m*k essentially telling us we have an identity

jimthompson5910 (jim_thompson5910):

So essentially, if you have a linear equation going through the origin (0,0), then it will satisfy the property that f(x+3) = f(x) + f(3)

OpenStudy (adamk):

@jim_thompson5910 I see that now. Thank you!

jimthompson5910 (jim_thompson5910):

no problem